Which plan has the least amount of risk?

Plan A
Payout P(Payout)
$25,000$ 0.26
$45,000 0.15
$60,000 0.59
Plan B
Payout P(Payout)
−$30,000 0.52
$15,000 0.2
$60,000 0.28
Answer #1

Expected amount of plan A = (0.26*25000)+(0.15*45000)+(0.59*60000) = 6500 + 6750 + 35400 = $48650

Expected amount of plan B = (-0.52*30000)+(0.2*15000)+(0.28*60000) = -15600 + 3000 + 16800 = $4200

So, Plan A has the least amount of risk